# Charged-Current Elastic Scattering at the Electron-Ion Collider

### v2

@@ I Introduction
Figure 1: Diagram of the CCE process. In the elastic process, unlike charged current deep inelastic scattering, the $W^{-}$ does not resolve the partonic structure of the nucleon and instead is sensitive to the axial form factor.
@@ V Projections
VI Remarks on Charged Current Cross Sections
It can be seen from Fig. 2 that the CCE total cross section remains approximately constant with $\sqrt{s}$ due to competition between the nucleon form factors and increased effective flux of $W^{-}$. On the other hand, the inclusive charged-current cross section increases linearly with $E_{\nu}$ due to increasing contribution of charged-current DIS, which dominates above $E_{\nu}\approx 20$ GeV. The phase space in $\gamma p$ center-of-mass energy $W$ for hard exclusive processes such as deeply virtual Compton scattering (DVCS) and deeply virtual meson production (DVMP) are much larger than for elastic scattering, and the DVCS and DVMP cross sections are known to grow with $W$ H1:2005gdw; ZEUS:2008hcd; H1:2009cml; Kumericki:2009uq. It may be the case then that at EIC energies, the cross sections for charged-current exclusive processes such as CCDVCS ($e^{-}p\rightarrow\nu_{e}n\gamma$) or CCDVMP (e.g. $e^{-}p\rightarrow\nu_{e}p\pi^{-}$) could be larger than the CCE cross section and counterintuitively may be more experimentally favorable. In particular, the $e^{-}p\rightarrow\nu_{e}p\pi^{-}$ reaction would have improved resolution on missing transverse momentum and missing mass since they could be measured in trackers. The charged-current pion production process is also expected to have large spin asymmetries that can be exploited Graczyk:2021oyl; Graczyk:2023lrm. Theoretical estimates of these cross sections at EIC energies would therefore be interesting to see.
@@ Acknowledgments
I would like to acknowledge Tim Hobbs, Garth Huber, Sylvester Joosten, Jihee Kim, Zein-Eddine Meziani, Noémie Pilleux, and Maria Żurek for useful discussions and comments. I also thank Lindsay DeWitt for editing this manuscript. This work was supported by the U.S. Department of Energy under Contract No. DE-AC02-06CH11357.

### v3

@@ I Introduction
Figure 1: Diagram of the CCE process. In the elastic process, unlike charged-current deep inelastic scattering, the $W^{-}$ does not resolve the partonic structure of the nucleon and instead is sensitive to the axial form factor.
@@ V Projections
VI Remarks on Other Charged-Current Cross Sections
It can be seen from Fig. 2 that the CCE total cross section remains approximately constant with $\sqrt{s}$ due to competition between the nucleon form factors and increased effective flux of $W^{-}$. On the other hand, the inclusive charged-current cross section increases linearly with $E_{\nu}$ due to increasing contribution of charged-current DIS, which dominates above $E_{\nu}\approx 20$ GeV. The phase space in $\gamma p$ center-of-mass energy $W$ for hard exclusive processes such as deeply virtual Compton scattering (DVCS) and deeply virtual meson production (DVMP) are much larger than for elastic scattering, and the DVCS and DVMP cross sections are known to grow with $W$ H1:2005gdw; ZEUS:2008hcd; H1:2009cml; Kumericki:2009uq. One may expect then that at EIC energies, the cross sections for charged-current exclusive processes such as CCDVCS ($e^{-}p\rightarrow\nu_{e}n\gamma$) or CCDVMP (e.g. $e^{-}p\rightarrow\nu_{e}p\pi^{-}$) could be larger than the CCE cross section and counterintuitively may be more experimentally favorable. This indeed appears to be the case, as the cross sections for $\pi^{-}$ and $D_{s}^{-}$ meson production at the EIC were calculated using GPD models to be reasonably large compared to the CCE cross section in Refs. Siddikov:2019ahb; Pire:2021dad. While the combination of relatively low cross section and unfavorable experimental branching fractions may make $D_{s}^{-}$ challenging, the CCDVMP $e^{-}p\rightarrow\nu_{e}p\pi^{-}$ reaction could be a nice experimental target for the EIC since the $\pi^{-}$ is measured directly and not through a decay. The all-charged hadronic final-state of $p\pi^{-}$ will also have excellent resolution on missing mass and missing transverse momentum since the particles can both be measured in trackers. The charged-current meson production process is also expected to have large spin asymmetries that can be exploited similarly to CCE scattering Pire:2021dad; Graczyk:2021oyl; Graczyk:2023lrm.
@@ Acknowledgments
I would like to acknowledge Tim Hobbs, Garth Huber, Sylvester Joosten, Jihee Kim, Zein-Eddine Meziani, Noémie Pilleux, and Maria Żurek for useful discussions and comments. I thank Bernard Pire for bringing Refs. Siddikov:2019ahb; Pire:2021dad to my attention. I also thank Lindsay DeWitt for editing this manuscript. This work was supported by the U.S. Department of Energy under Contract No. DE-AC02-06CH11357.